Kyky232:

A circle is centered at (7, 8) and has a radius of 11. Which of the following is the equation for this circle? (x − 7)2 + (y − 8)2 = 121 (x − 7)2 + (y − 8)2 = 11 (x + 7)2 + (y + 8)2 = 121 (x + 7)2 + (y + 8)2 = 11

AverageStudent:

 what are need to solve for? \[(a-x)^2+(b-y)^2=r^2\]

Kyky232:

~i dont know~ thats why i asked the question

AverageStudent:

so, none of numbers are negative; \[(a-7)^2+(b-8)^2=11\times11=121\] So; \[(a-7)^2+(b-8)^2=121\]

AverageStudent:

please don't get mad, Im trying to help

Kyky232:

@averagestudent wrote:
please don't get mad, Im trying to help
im not getting mad im just confused with the question lol. Geometry really dont stick with me

AverageStudent:

Oh ok. I solve it for you.

AverageStudent:

up there

AverageStudent:

@averagestudent wrote:
so, none of numbers are negative; \[(a-7)^2+(b-8)^2=11\times11=121\] So; \[(a-7)^2+(b-8)^2=121\]

AverageStudent:

a=x b=y

Kyky232:

so A?

AverageStudent:

Yes. Congrats!